An internally reversible Rankine cycle with both reheat and regeneration operates between 400OC and 3.8 MPa and 50OC. The first stage turbine exhausts 700 kPa where part part of the steam is extracted for use in an open feed-water heater. The remainder is heated to 500 OC before entering the second stage. Saturated steam leaves the feedwater heater. Compute the cycle efficiency. USE THE T-s CHART TO FIND THE ENTHALPIES. Neglect pump work.
Answer #1

\textbf{Given data}\rightarrow

P1 3.8 MPa

T_{1}=400\: \: \: ^{o}\textup{C}

P_{2}=700\textup{ kPa}

P3 = 700 kPa

T_{3}=500\: \: \: ^{o}\textup{C}

P_{4}=700\textup{ kPa}

T_{5}=50\: \: \: ^{o}\textup{C}

T-s diagram →

1596239391012_image.png

\textbf{Solution}\rightarrow

\textup{from T-s diagram enthalpy calculated from various points are }

h_{1}=3217.86\textup{ kJ/kg}

s_{1}=6.799\textup{ kJ/kgK}

s_{1}=s_{2}

h_{2}=2803.06\textup{ kJ/kg}

h_{3}=3481.81\textup{ kJ/kg}

s_{3}=7.932\textup{ kJ/kgK}

h_{4}=2803.06\textup{ kJ/kg}

s_{3}=s_{5}

h_{5}=2545.061\textup{ kJ/kg}

h_{7}=h_{f}\textup{ at 50 }^{o}\textup{C}=209.34\textup{ kJ/kg}

h_{11}=h_{f}\textup{ at 700 kPa }=697\textup{ kJ/kg}

\textbf{Solution}\rightarrow

\textup{energy balance for open feed water heater}

\left ( 1-y \right )h_{7}+y*\left ( h_{4} \right )=1*h_{11}

\left ( 1-y \right )*209.31+y*2803.06=1*697

209.34-209.34y+y*2803.06=1*697

y=\frac{697-209.34}{2803.06-209.34}

y=0.188

Turbine work (WT) →

wt = (hi – h2) + (1 - y) * (h2 – h4) + (1 - y) * (h3 - 95)

w_{T}=\left ( 3217.86-2803.06 \right )+\left ( 1-0.188 \right )*\left ( 2803.06-2803.06 \right )+\left ( 1-0.188 \right )*\left ( 3481.81-2545.061 \right )

w_{T}=1175.44\textup{ kJ/kg}

Heat supplied in boiler (3) ►

q_{s}=\left ( h_{1}-h_{11} \right )+\left ( 1-y \right )*\left ( h_{3}-h_{2} \right )

q_{s}=\left ( 3217.86-697 \right )+\left ( 1-0.188 \right )*\left ( 3481.81-2803.06 \right )

q_{s}=3072.005\textup{ kJ/kg}

\eta =\frac{w_{T}}{q_{s}}

\eta =\frac{1175.44}{3072.005}

\eta =0.3826